OpenStudy (anonymous):

I have a problem where they want me to find the radius of a cone with a volume of 60 and a height of 21?

OpenStudy (jdoe0001):

\(\bf \textit{volume of a cone}=\cfrac{\pi {\color{blue}{ r}}^2 h}{3}\qquad h=height=21\qquad volume=60 \\ \quad \\ 60=\cfrac{\pi {\color{blue}{ r}}^2 \cdot 21}{3}\) solve for "r", or radius